OpenStudy (anonymous):

how to take the derivative of e^x^2

OpenStudy (abb0t):

chain rule. derivative of the outside times the derivative of the inside.

OpenStudy (anonymous):

Is it: \[\large e^{x^2}\]

OpenStudy (abb0t):

\[\frac{ d }{ db }e^b = \frac{ d }{ db }(e^b) \times \frac{ d }{ db }(b)=\]

OpenStudy (abb0t):

Not quite, @waterineyes :) You forgot to take the derivative of "b" like in the example I provided. You can think of it as: \(b=x^2\)

OpenStudy (anonymous):

I just confirmed the question from asker..

OpenStudy (anonymous):

so it will be 2e^x multiplied by e^x multiplied by 1

OpenStudy (anonymous):

@abb0t so it will be 2e^x multiplied by e^x multiplied by 1

OpenStudy (abb0t):

My mistake @waterineyes No, @0213 It is just the derivative of \(e^x\) but in this case, you have \(e^{x^2}\) so it will stay the same, multiplied by the derivative of \(x^2\) :)

OpenStudy (anonymous):

so e^x^2 multiplied by 2x

OpenStudy (anonymous):

@abb0t so e^x^2 multiplied by 2x

OpenStudy (abb0t):

Correct :)

OpenStudy (abb0t):

Derivative of the OUTSIDE times derivative of the INSIDE \(e^x\) is the "outside" while \(x^2\) is the inside, because it's inside of another function, in this case, the exponent.

OpenStudy (anonymous):

could you answer another question

OpenStudy (abb0t):

Sure.

OpenStudy (anonymous):

how

OpenStudy (abb0t):

What do you mean how would I have?

OpenStudy (anonymous):

sorry, my key bored is not working right. I mean how would I

OpenStudy (anonymous):

how would i find the x-intercepts of 2xe^x^2 @abb0t

OpenStudy (abb0t):

Set it equal to zero and solve for "x" if possible.

OpenStudy (anonymous):

is there none because of e and the Ln of 0 is undefined

OpenStudy (abb0t):

What about x = 0 :) Yes?

OpenStudy (anonymous):

would I have to solve for 2x and e^x^2 as one function or two?

OpenStudy (abb0t):

One. You can divide \(e^{x^2}\) by zero. to get 2x=0 and then get x =0